May 21, 2013 Calorimetry

From GlueXWiki
Jump to: navigation, search

'Teleconference Time: 11:00 a.m. EST

  • ESNET (Number is 8542553) and EVO session (GlueX Calorimetry meeting room)
  • Phone connection only upon request.
    • +1-866-740-1260 : US+Canada
    • +1-303-248-0285 : International
    • then enter participant code: 3421244# (remember the "#").
    • or (and code without the #)

Participant Direct Lines

  • JLab Phone: in CC F326 is 757-269-6460 (usual room)
  • JLab Phone in CC L207 is 757-269-7084
  • Phone in the Regina Videoconference Suite is 306-585-4204
  • Athens Phone: in Christina's office is 011-30-210-727-6947

Items for followup from previous meeting(s)

  1. Light guides
    • Return 84 non-conforming light guides (Elton)
    • Prepare letter for customs to prepare for exchange (Alison)

Tentative Agenda

  1. Announcements
  2. Collaboration Meeting Schedule
  3. Action Items
  4. Monitoring
  5. FCAL
  6. BCAL
  7. Reconstruction and Simulation
  8. TDR update
  9. Calibration of Calorimeters
  10. Any other business


Attendees: David, Beni, Yi (JLab), Will (USM), Christina, George Vasileiadis, Pavlos (Athens), Shaun, Andrei, Zisis (Regina)

  1. Announcements: Christina announced that the FCAL boards plus 3 SiPMs plus plexiglas have been shipped to JLab and should arrive tomorrow.
  2. Collaboration Meeting Schedule was finalized
  3. Action Items: defer until Elton returns
  4. Monitoring: No other report. Christina requested that Matt approve the FCAL Monitoring manual (which he did after the meeting). Beni and Zisis approved the BCAL one. Beni will send an email to this effect so that Athens can be paid.
  5. FCAL: Matt reported that Dan Bennett is at IU working on the remaining basis. We have 3001 and another 15 that will stay in Indiana. This satisfies the contract. Congrats to IU. Dan will go to JLab soon to finish the remaining blocks for the arrau. Dark room is progressing.
  6. BCAL:
    1. Zisis reported that the old (Mehrnoosh) PDE data was unreliable. Two large, offsetting errors were found, that fortuitously cancelled each other so that the PDE came out roughly ok. The apparatuis has been modified to fix the issues and the methodology corrected. All measurements have been repeated by Tegan from the ground up and these will be published. The paper will be circulated to the collaboration soon.
    2. Shaun is progressing on his cuts on the cosmic runs at JLab, aiming to extract the energy deposition in each cell.
  7. Tegan is continuing the cold measurements for the extraction of PDE and dark contribution.
  8. Reconstruction and Simulation: nothing to report.
  9. TDR update: George Lolos has give comments to Zisis on the BCAL section, which will be incorporated before the collab meeting and then opened up to all for commenting.
  10. Calibration of Calorimeters: nothing to report
  11. Any other business: none